OpenStudy (stuck-help):

METAL square root 5q^2-20=0

OpenStudy (bohotness):

Home | Latest | Discuss | About | Help | Translation 8,197,024 solved | 566 online -5q+20=3q2 Two solutions were found : q =(-5-√265)/6=-3.546 q =(-5+√265)/6= 1.880 Rearrange: Rearrange the equation by subtracting what is to the right of the equal sign from both sides of the equation : -5*q+20-(3*q^2)=0 Step by step solution : Step 1 : Raise q to the 2nd power Exponentiaion : Equation at the end of step 1 : (20 - 5q) - (3 • q2) = 0 Step 2 : Simplify 20-5q - 3q2 Pulling out like terms : 2.1 Pull out like factors : -3q2 - 5q + 20 = -1 • (3q2 + 5q - 20) Trying to factor by splitting the middle term 2.2 Factoring 3q2 + 5q - 20 The first term is, 3q2 its coefficient is 3 . The middle term is, +5q its coefficient is 5 . The last term, "the constant", is -20 Step-1 : Multiply the coefficient of the first term by the constant 3 • -20 = -60 Step-2 : Find two factors of -60 whose sum equals the coefficient of the middle term, which is 5 . -60 + 1 = -59 -30 + 2 = -28 -20 + 3 = -17 -15 + 4 = -11 -12 + 5 = -7 -10 + 6 = -4 -6 + 10 = 4 -5 + 12 = 7 -4 + 15 = 11 -3 + 20 = 17 -2 + 30 = 28 -1 + 60 = 59 Observation : No two such factors can be found !! Conclusion : Trinomial can not be factored Equation at the end of step 2 : -1 • (3q2 + 5q - 20) = 0 Step 3 : Solve -1•(3q2+5q-20) = 0 Equations which are never true : 3.1 Solve : -1 = 0 This equation has no solution. A a non-zero constant never equals zero. Parabola, Finding the Vertex : 3.2 Find the Vertex of y = 3q2+5q-20 Parabolas have a highest or a lowest point called the Vertex . Our parabola opens up and accordingly has a lowest point (AKA absolute minimum) . We know this even before plotting "y" because the coefficient of the first term, 3 , is positive (greater than zero). Each parabola has a vertical line of symmetry that passes through its vertex. Because of this symmetry, the line of symmetry would, for example, pass through the midpoint of the two x -intercepts (roots or solutions) of the parabola. That is, if the parabola has indeed two real solutions. Parabolas can model many real life situations, such as the height above ground, of an object thrown upward, after some period of time. The vertex of the parabola can provide us with information, such as the maximum height that object, thrown upwards, can reach. For this reason we want to be able to find the coordinates of the vertex. For any parabola,Aq2+Bq+C,the q -coordinate of the vertex is given by -B/(2A) . In our case the q coordinate is -0.8333 Plugging into the parabola formula -0.8333 for q we can calculate the y -coordinate : y = 3.0 * -0.83 * -0.83 + 5.0 * -0.83 - 20.0 or y = -22.083 Parabola, Graphing Vertex and X-Intercepts : Root plot for : y = 3q2+5q-20 Axis of Symmetry (dashed) {q}={-0.83} Vertex at {q,y} = {-0.83,-22.08} q -Intercepts (Roots) : Root 1 at {q,y} = {-3.55, 0.00} Root 2 at {q,y} = { 1.88, 0.00} Solve Quadratic Equation by Completing The Square 3.3 Solving 3q2+5q-20 = 0 by Completing The Square . Divide both sides of the equation by 3 to have 1 as the coefficient of the first term : q2+(5/3)q-(20/3) = 0 Add 20/3 to both side of the equation : q2+(5/3)q = 20/3 Now the clever bit: Take the coefficient of q , which is 5/3 , divide by two, giving 5/6 , and finally square it giving 25/36 Add 25/36 to both sides of the equation : On the right hand side we have : 20/3 + 25/36 The common denominator of the two fractions is 36 Adding (240/36)+(25/36) gives 265/36 So adding to both sides we finally get : q2+(5/3)q+(25/36) = 265/36 Adding 25/36 has completed the left hand side into a perfect square : q2+(5/3)q+(25/36) = (q+(5/6)) • (q+(5/6)) = (q+(5/6))2 Things which are equal to the same thing are also equal to one another. Since q2+(5/3)q+(25/36) = 265/36 and q2+(5/3)q+(25/36) = (q+(5/6))2 then, according to the law of transitivity, (q+(5/6))2 = 265/36 We'll refer to this Equation as Eq. #3.3.1 The Square Root Principle says that When two things are equal, their square roots are equal. Note that the square root of (q+(5/6))2 is (q+(5/6))2/2 = (q+(5/6))1 = q+(5/6) Now, applying the Square Root Principle to Eq. #3.3.1 we get: q+(5/6) = √ 265/36 Subtract 5/6 from both sides to obtain: q = -5/6 + √ 265/36 Since a square root has two values, one positive and the other negative q2 + (5/3)q - (20/3) = 0 has two solutions: q = -5/6 + √ 265/36 or q = -5/6 - √ 265/36 Note that √ 265/36 can be written as √ 265 / √ 36 which is √ 265 / 6 Solve Quadratic Equation using the Quadratic Formula 3.4 Solving 3q2+5q-20 = 0 by the Quadratic Formula . According to the Quadratic Formula, q , the solution for Aq2+Bq+C = 0 , where A, B and C are numbers, often called coefficients, is given by : - B ± √ B2-4AC q = ———————— 2A In our case, A = 3 B = 5 C = -20 Accordingly, B2 - 4AC = 25 - (-240) = 265 Applying the quadratic formula : -5 ± √ 265 q = —————— 6 √ 265 , rounded to 4 decimal digits, is 16.2788 So now we are looking at: q = ( -5 ± 16.279 ) / 6 Two real solutions: q =(-5+√265)/6= 1.880 or: q =(-5-√265)/6=-3.546 Two solutions were found : q =(-5-√265)/6=-3.546 q =(-5+√265)/6= 1.880